You have a mass of 60kg. Your car, traveling at 10m/s, goes down a dip in the road with a radius of curvature of 60m. How heavy do you feel? N
Answer #1

At the lowest point of the dip, the centripetal force and weight will be acting vertically downwards.

Apparent weight, W_{app}=W+F_{c}

W_{app}=mg+\frac{mv^{2}}{r}

W_{app}=60kg*9.81m/s^{2}+\frac{60kg*(10m/s)^{2}}{60m}

W_{app}=60*9.81+\frac{60*10^{2}}{60}

W_{app}=588.6+100

ANSWER: {\color{Red} W_{app}=688.6N}

==========================